### v1

# Ordnung muss sein

Abstract.
For any length category, we establish a set of rules (necessary and sufficient) that ensure a partial order on the isomorphism classes of simple objects such that the category is equivalent to the category of finite dimensional representations of this partially ordered set.
Key words and phrases:
Length category, ext-quiver, representation, partially ordered set
2020 Mathematics Subject Classification:
16G20 (primary); 18E10 (secondary)
1. Introduction
A length category is an abelian category such that any chain of subobjects for a given object is finite and the isomorphism classes of objects form a set [6]. A basic example is given by the category of finite dimensional representations of a finite poset. Here, we work over a fixed field and the following result provides a characterisation of the length categories that arise in this way. [^1] Following the title of the paper, we may refer to conditions (1)–(3) as ‘Hausordnung’, cf. https://de.wikipedia.org/wiki/Hausmeister_Krause_-_Ordnung_muss_sein.
Theorem 1.
Let $\mathbb{K}$ be a field. A $\mathbb{K}$-linear length category is equivalent to the category of finite dimensional representations of a finite poset if and only if the following holds:
  1. (1)
    The ext-quiver has no oriented cycle.
  2. (2)
    The endomorphism ring of each simple object is isomorphic to $\mathbb{K}$.
  3. (3)
    There exists an object $M$ with a distributive lattice of subobjects such that each object is a subquotient of a finite direct sum of copies of $M$.
In this case the poset is given by the isomorphism classes of simple objects, with $S]\leq[T]$ if there is a path $S]\to\cdots\to[T]$ in the ext-quiver. In particular, the Hasse diagram of the poset identifies with the ext-quiver.
A fundamental discrete invariant of a length category is its ext-quiver. The vertices are given by the isomorphism classes of simple objects, and there is an arrow $S]\to[T]$ if $\operatorname{Ext}^{1}(T,S)\neq 0$. The above theorem describes the length categories that are completely determined by their ext-quiver.
There is a long tradition of studying representations of posets, and we refer to the monograph of Simson for a detailed account [9]. Intimately related is the study of incidence algebras and their modules. In fact, the question when a finite dimensional algebra is an incidence algebra has been investigated by several authors [1, 5, 8]. In particular, the relevance of distributive modules has been noticed, but the more general context of length categories seems to be new.
2. The proof
Distributive objects
Fix a length category. For any object $M$ and a simple object $S$ let $[M:S]$ denote the multiplicity of $S$ in a composition series of $M$. The object $M$ is multiplicity free if $M:S]\leq 1$ for every simple object $S$.
Recall that an object has a distributive lattice of subobjects if and only if there is no simple object $S$ such that $S\oplus S$ arises as a subquotient; see for instance the first exercise in [2]. Objects with this property are called distributive.
Lemma 2.
Every multiplicity free object is distributive. The converse holds when the ext-quiver has no oriented cycles.
Proof.
The first assertion is clear, given the characterisation of distributive objects. Now fix an object $M$ and set $M^{i}:=\operatorname{rad}^{i}M$ for $i\geq 0$. Then $M^{r}=0$ for some $r\geq 0$, and this yields a filtration
$0=M^{r}\subseteq\cdots\subseteq M^{1}\subseteq M^{0}=M$
with semisimple factors. For any pair of simple objects $S\subseteq M^{i}/M^{i+1}$ and $T\subseteq M^{i-1}/M^{i}$ there exists an arrow $S]\to[T]$ in the ext-quiver, which comes from the extension
$0\longrightarrow M^{i}/M^{i+1}\longrightarrow M^{i-1}/M^{i+1}\longrightarrow M^{i-1}/M^{i}\longrightarrow 0.$
If $[M:S]>1$ for a simple object $S$, then $S$ arises in $M^{i}/M^{i+1}$ and $M^{j}/M^{j+1}$ for some pair $i,j$. When $M$ is distributive, then $i\neq j$. Thus there exists a cycle in the ext-quiver. ∎
Representations of posets
Fix a finite poset $P$ and a field $\mathbb{K}$. It is convenient to view $P$ as a category: the objects are the elements of $P$ and there is a unique morphism $x\to y$ if and only if $x\leq y$. We write $\operatorname{rep}(P,\mathbb{K})$ for the category of contravariant representations, so
$\operatorname{rep}(P,\mathbb{K}):=\operatorname{Fun}(P^{\mathrm{op}},\operatorname{mod}\mathbb{K}).$
For any ring $A$ let $\operatorname{mod}A$ denote the category of finitely presented right $A$-modules. We write $\mathbb{K}P$ for the incidence algebra which is a $\mathbb{K}$-linear space with basis given by elements $a_{xy}$ for any pair $x\leq y$ in $P$ and multiplication induced by
$a_{xy}a_{vw}:=\begin{cases}a_{vy}&\text{if }w=x,\\ 0&\text{otherwise}.\end{cases}$
For $M$ in $\operatorname{rep}(P,\mathbb{K})$ there is a canonical right action of $\mathbb{K}P$ on $\bigoplus_{x\in P}M(x)$.
Lemma 3.
The assignment $M\mapsto\bigoplus_{x\in P}M(x)$ induces an equivalence
$\operatorname{rep}(P,\mathbb{K})\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}\mathbb{K}P.$
Proof.
Taking a $\mathbb{K}P$-module $N$ to the representation of $P$ given by
$x\mapsto N(x):=Na_{xx}$
and for $x\leq y$ the linear map $N(y)\to N(x)$ given by right multiplication with $a_{xy}$ yields a quasi-inverse. ∎
Representations of algebras
We recall a few basic facts about modules over finite dimensional algebras. Let $A$ be a right artinian ring and fix a decomposition $1=\sum_{x\in P}e_{x}$ into pairwise orthogonal and primitive idempotents. We may assume that $A$ is basic. Thus the indecomposable projective modules are of the form $e_{x}A$ for some index $x$, and
$\operatorname{Hom}_{A}(e_{x}A,M)=Me_{x}$
for any $A$-module $M$. Let $J=J(A)$ denote the Jacobson radical of $A$. Then the simple modules are of the form $S_{x}:=e_{x}(A/J)$ for some index $x$. Set $E_{x}:=\operatorname{End}_{A}(S_{x})$.
The following lemma is the basic tool that relates the ext-quiver of $\operatorname{mod}A$ to the structure of the algebra $A$.
Lemma 4.
There is an an isomorphism of $E_{y}$-$E_{x}$-bimodules
$\operatorname{Ext}_{A}^{1}(S_{x},S_{y})\cong\operatorname{Hom}_{A}(e_{x}(J/J^{2}),S_{y})\cong\operatorname{Hom}_{E_{y}}(e_{x}(J/J^{2})e_{y},E_{y}).$
Proof.
See for example Proposition 2.4.3 in [2]. ∎
Now assume that the ext-quiver of $\operatorname{mod}A$ is acyclic. Then $\operatorname{End}_{A}(e_{x}A)\cong E_{x}$ for each $x$ by the above lemma. For any $A$-module $M$ of finite length, an induction on its length shows that the multiplicity of $S_{x}$ in a composition series is given by
$M:S_{x}]=\operatorname{dim}_{E_{x}}\operatorname{Hom}_{A}(e_{x}A,M).$
We assume in addition that $A$ is a $\mathbb{K}$-algebra and that $E_{x}\cong\mathbb{K}$ for all $x$.
Lemma 5.
Let $M$ be an $A$-module such that $M:S_{x}]=1$ and $e_{x}A$ is a subquotient of $M^{r}$ for some $r\geq 1$. Then every non-zero morphism $e_{x}A\to M$ is a monomorphism.
Proof.
The assumption yields a monomorphism $\phi\colon e_{x}A\to M^{r}$ for some $r\geq 1$. Choose a non-zero morphism $\psi\colon e_{x}A\to M$. Then there are $\lambda_{i}\in E_{x}$ such that $\phi_{i}=\psi\lambda_{i}$ for $1\leq i\leq r$. We have
$\operatorname{Ker}\psi=\bigcap_{i}\operatorname{Ker}\psi\lambda_{i}=\operatorname{Ker}\phi=0$
where the first equality uses that the $\lambda_{i}$ are central. ∎
Now fix an $A$-module $M$ such that $M:S_{x}]=1$ for all $x$, and we assume that every non-zero morphism $e_{x}A\to M$ is a monomorphism. Then $e_{x}A$ is multiplicity free for all $x$ and there is a $\mathbb{K}$-basis of $A$ that consists of elements $e_{xy}\in e_{y}Ae_{x}$. In particular, there is a partial order on the set $P$ given by
$x\leq y\quad:\iff\quad e_{y}Ae_{x}\neq 0,$
and for $x\leq y\leq z$ there is an element $\varepsilon_{xz}^{y}\in\mathbb{K}$ such that
$e_{yz}e_{xy}=e_{xz}\varepsilon_{xz}^{y}.$
Next we choose a $\mathbb{K}$-basis of $M$ that consists of elements $m_{x}\in Me_{x}$. Then there are non-zero elements $\mu_{xy}\in\mathbb{K}$ for all $x\leq y$ such that
$m_{y}e_{xy}=m_{x}\mu_{xy}.$
The equality
$m_{z}(e_{yz}e_{xy})=(m_{z}e_{yz})e_{xy}$
gives
($\ast$) $\mu_{xz}\varepsilon_{xz}^{y}=\mu_{yz}\mu_{xy}.$
We are ready to compare $A$ with the incidence algebra $\mathbb{K}P$ that is given by the poset coming from the ext-quiver of $\operatorname{mod}A$.
Lemma 6.
The $\mathbb{K}$-linear map $\phi\colon A\to\mathbb{K}P$ given by the assignment $e_{xy}\mapsto a_{xy}\mu_{xy}$ for any $x\leq y$ is an isomorphism of $\mathbb{K}$-algebras.
Proof.
The map is an isomorphism of vector spaces and we need to show that it is multiplicative. For $x\leq y\leq z$ we compute
$\phi(e_{yz}e_{xy})$ $=\phi(e_{yz}\varepsilon_{xz}^{y})$
$=a_{xz}\mu_{xz}\varepsilon_{xz}^{y}$
[…]
$=(a_{yz}a_{xy})\mu_{yz}\mu_{xy}$
$=(a_{yz}\mu_{yz})(a_{xy}\mu_{xy})$
$=\phi(e_{yz})\phi(e_{xy})$
and this yields the assertion. ∎
Proof of the main theorem
Recall that the Hasse diagram of a poset $P$ is the quiver with vertex set $P$ and an arrow $x\to y$ if $x<y$ and there is no $z$ with $x<z<y$.
Proof of Theorem 1.
Let $\mathcal{A}$ be a $\mathbb{K}$-linear length category. Suppose first that $\mathcal{A}$ is equivalent to $\operatorname{rep}(P,\mathbb{K})$ for some finite poset $P$. The simple representations identify with the elements of $P$. Specifically, for $x\in P$ the simple representation $S_{x}$ is given by $S_{x}(y)=\delta_{xy}\mathbb{K}$ for $y\in P$. In particular, it satisfies $\operatorname{End}(S_{x})\cong\mathbb{K}$. The ext-quiver identifies with the Hasse diagram of $P$. Thus there are no oriented cycles. Now consider the representation $M$ given by $M(x)=\mathbb{K}$ and $M(y)\to M(x)$ the identity map for all $x\leq y$. Then $M:S_{x}]=1$ for all $x$, and therefore the lattice of subobjects is distributive. The subrepresentation $M_{x}\subseteq M$ satisfying $M_{x}(y)\neq 0$ if and only if $y\leq x$ yields a projective cover of $S_{x}$. Thus every indecomposable projective representation embeds into $M$, and each object is a subquotient of $M^{r}$ for some $r\geq 1$.
Now suppose that conditions (1)–(3) hold for $\mathcal{A}$. It follows from [3, Proposition 2.14] that $\mathcal{A}$ is equivalent to the category $\operatorname{mod}A$ for some finite dimensional $\mathbb{K}$-algebra $A$. More precisely, conditions (2)–(3) imply that the category $\mathcal{A}$ admits a projective generator $Q$ such that $\operatorname{Hom}(Q,-)$ induces an equivalence $\mathcal{A}\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}A$ for $A=\operatorname{End}(Q)$. Using Lemmas 2 and 5, the existence of a distributive object implies that Lemma 6 can be applied. Thus $A\cong\mathbb{K}P$ for the poset $P$ given by the isomorphism classes of simple $A$-modules, and the equivalence with $\operatorname{rep}(P,\mathbb{K})$ follows from Lemma 3. ∎
Remark 7.
The distributive object $M$ arising in condition (3) of Theorem 1 is not necessarily unique. In fact, such objects are parameterised by 22-cocycles $\mu=(\mu_{xy})$ in $H^{*}(\Sigma(P),\mathbb{K}^{\times})$, where $\Sigma(P)$ denotes the simplicial complex associated with $P$; see [7, 8].
Remark 8.
Condition (1) of Theorem 1 is needed. The category of finite length modules over the algebra $\mathbb{K}[\varepsilon]$ of duals numbers yields an example where (2) and (3) hold. Take for $M$ the indecomposable projective module.
The following example is interesting for several reasons, and I am grateful to Claus Ringel for suggesting it. The example shows that condition (2) of Theorem 1 is needed. Also, it yields infinitely many objects $M$ of length two such that all objects are a subquotient of a finite direct sum of copies of $M$.
Example 9.
Consider the matrix algebra
$A=\begin{bmatrix}\mathbb{C}&\mathbb{C}\oplus\mathbb{C}_{\sigma}\\ 0&\mathbb{C}\end{bmatrix}$
where the action of $\mathbb{C}$ on $\mathbb{C}_{\sigma}$ from the right is twisted by conjugation; it is a finite dimensional $\mathbb{R}$-algebra. There are infinitely many isomorphism classes of length two $A$-modules, and all but two are faithful; see the Addendum of [4].
Remark 10.
It would be interesting to have a version of Theorem 1 characterising categories of the form $\operatorname{rep}(P,\mathbb{K})$ when $\mathbb{K}$ is a skew field.
Acknowledgements
The author is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support. This work was supported by the Deutsche Forschungsgemeinschaft (SFB-TRR 358/1 2023 - 491392403).