### v2

# Ordnung muss sein

Abstract.
For any length category, we establish a set of rules (necessary and sufficient) that ensure a partial order on the isomorphism classes of simple objects such that the category is equivalent to the category of finite dimensional representations of this partially ordered set.
Key words and phrases:
Length category, ext-quiver, centre, partially ordered set, linear representation, incidence algebra, distributive object
2020 Mathematics Subject Classification:
16G20 (primary); 18E10 (secondary)
1. Introduction
A length category is an abelian category such that any chain of subobjects for a given object is finite and the isomorphism classes of objects form a set [8]. There are two basic invariants: the centre, which is the ring of natural transformations from the identity functor to itself, and the ext-quiver, for which the vertices are given by the isomorphism classes of simple objects, and there is an arrow $S]\to[T]$ if $\operatorname{Ext}^{1}(T,S)\neq 0$. The following result describes a class of length categories that are completely determined by these invariants. [^1] Following the title of the paper, we may refer to conditions (1)–(3) as ‘Hausordnung’, cf. https://de.wikipedia.org/wiki/Hausmeister_Krause_-_Ordnung_muss_sein.
For an abelian category $\mathcal{A}$ we write $Z(\mathcal{A})$ for its centre and say that an endomorphism of an object $X$ is central if it lies in the image of the canonical map $Z(\mathcal{A})\to\operatorname{End}(X)$. The category $\mathcal{A}$ is connected if 0 and $1$ are the only idempotents in $Z(\mathcal{A})$.
Theorem 1.
A connected length category is equivalent to the category of finite dimensional representations of a finite poset over a field if and only if the following hold:
  1. (1)
    The ext-quiver has no oriented cycle.
  2. (2)
    For each simple object all its endomorphisms are central.
  3. (3)
    There exists an object $M$ with a distributive lattice of subobjects such that each object is a subquotient of a finite direct sum of copies of $M$.
In this case the field equals the centre of the category and the poset is given by the isomorphism classes of simple objects, with $S]\leq[T]$ if there is a path from $[S]$ to $[T]$ in the ext-quiver.
There is a long tradition of studying linear representations of posets, and we refer to the monograph of Simson for a detailed account [11]. Intimately related is the study of incidence algebras and their modules. In fact, the question when a finite dimensional algebra is an incidence algebra has been investigated by several authors [2, 6, 10]. In particular, the relevance of distributive modules has been noticed, but the more general context of length categories seems to be new.
Categories of linear representations of finite posets over a field are nothing but categories of sheaves on finite $T_{0}$-spaces with values in the category of vector spaces; see Remark 14. This yields another perspective on the class of length categories arising in Theorem 1, which is reminiscent of Gabriel’s reconstruction of a noetherian scheme from its category of quasi-coherent sheaves [7, VI].
2. The proof
Distributive objects
Fix a length category. For any object $M$ and a simple object $S$ let $[M:S]$ denote the multiplicity of $S$ in a composition series of $M$. The object $M$ is multiplicity free if $M:S]\leq 1$ for every simple object $S$.
Recall that an object has a distributive lattice of subobjects if and only if there is no simple object $S$ such that $S\oplus S$ arises as a subquotient; see for instance the first exercise in [3]. Objects with this property are called distributive.
Lemma 2.
Every multiplicity free object is distributive. The converse holds when the ext-quiver has no oriented cycles.
Proof.
The first assertion is clear, given the characterisation of distributive objects. Now fix an object $M$ and set $M^{i}:=\operatorname{rad}^{i}M$ for $i\geq 0$. Then $M^{r}=0$ for some $r\geq 0$, and this yields a filtration
$0=M^{r}\subseteq\cdots\subseteq M^{1}\subseteq M^{0}=M$
with semisimple factors. For any pair of simple objects $S\subseteq M^{i}/M^{i+1}$ and $T\subseteq M^{i-1}/M^{i}$ there exists an arrow $S]\to[T]$ in the ext-quiver, which comes from the extension
$0\longrightarrow M^{i}/M^{i+1}\longrightarrow M^{i-1}/M^{i+1}\longrightarrow M^{i-1}/M^{i}\longrightarrow 0.$
If $[M:S]>1$ for a simple object $S$, then $S$ arises in $M^{i}/M^{i+1}$ and $M^{j}/M^{j+1}$ for some pair $i,j$. When $M$ is distributive, then $i\neq j$. Thus there exists a cycle in the ext-quiver. ∎
Module categories
We fix a length category $\mathcal{A}$ and recall from [4] a simple criterion for $\mathcal{A}$ to be a category of finite length modules. Let $\mathbb{K}$ be a commutative ring and suppose that $\mathcal{A}$ is $\mathbb{K}$-linear, so the morphism sets are $\mathbb{K}$-modules and the composition maps are bilinear. Then $\mathcal{A}$ is called hom-finite if $\operatorname{Hom}(X,Y)$ is a $\mathbb{K}$-module of finite length for each pair of objects $X,Y$.
For any ring $A$ let $\operatorname{mod}A$ denote the category of finitely presented right $A$-modules.
Proposition 3.
A hom-finite length category is equivalent to $\operatorname{mod}A$ for some Artin algebra $A$ if and only if there exists an object $M$ such that each object is a subquotient of a finite direct sum of copies of $M$.
Proof.
See Proposition 2.14 in [4]. ∎
Example 4.
Let $\mathcal{A}$ be a length category and view it as $\mathbb{K}$-linear category for $\mathbb{K}=Z(\mathcal{A})$. Suppose that for each simple object all its endomorphisms are central. Then $\mathcal{A}$ is hom-finite. More precisely, for each pair of objects $X,Y$ in $\mathcal{A}$ we have
$\ell_{\mathbb{K}}(\operatorname{Hom}(X,Y))\leq\ell(X)\cdot\ell(Y).$
Representations of posets
Fix a finite poset $P$ and a field $\mathbb{K}$. It is convenient to view $P$ as a category: the objects are the elements of $P$ and there is a unique morphism $x\to y$ if and only if $x\leq y$. We write $\operatorname{rep}(P,\mathbb{K})$ for the category of contravariant representations, so
$\operatorname{rep}(P,\mathbb{K}):=\operatorname{Fun}(P^{\mathrm{op}},\operatorname{mod}\mathbb{K}).$
Let $\mathbb{K}P$ denote the incidence algebra which is a $\mathbb{K}$-linear space with basis given by elements $a_{xy}$ for any pair $x\leq y$ in $P$ and multiplication induced by
(1) $a_{xy}a_{vw}:=\delta_{wx}a_{vy}.$
For $M$ in $\operatorname{rep}(P,\mathbb{K})$ there is a canonical right action of $\mathbb{K}P$ on $\bigoplus_{x\in P}M(x)$.
Lemma 5.
The assignment $M\mapsto\bigoplus_{x\in P}M(x)$ induces an equivalence
$\operatorname{rep}(P,\mathbb{K})\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}\mathbb{K}P.$
Proof.
Taking a $\mathbb{K}P$-module $N$ to the representation of $P$ given by
$x\mapsto N(x):=Na_{xx}$
and for $x\leq y$ the linear map $N(y)\to N(x)$ given by right multiplication with $a_{xy}$ yields a quasi-inverse. ∎
Representations of algebras
We recall a few basic facts about modules over finite dimensional algebras. Let $A$ be a right Artinian ring and fix a decomposition $1=\sum_{x\in P}e_{x}$ into pairwise orthogonal and primitive idempotents. We may assume that $A$ is basic and connected. Thus the indecomposable projective modules are of the form $e_{x}A$ for some index $x$, and
$\operatorname{Hom}_{A}(e_{x}A,M)=Me_{x}$
for any $A$-module $M$. Let $J=J(A)$ denote the Jacobson radical of $A$. Then the simple modules are of the form $S_{x}:=e_{x}(A/J)$ for some index $x$. Set
$E_{x}:=\operatorname{End}_{A}(S_{x})\cong e_{x}(A/J)e_{x}.$
The following lemma is the basic tool that relates the ext-quiver of $\operatorname{mod}A$ to the structure of the algebra $A$.
Lemma 6.
There is an an isomorphism of $E_{y}$-$E_{x}$-bimodules
$\operatorname{Ext}_{A}^{1}(S_{x},S_{y})\cong\operatorname{Hom}_{A}(e_{x}(J/J^{2}),S_{y})\cong\operatorname{Hom}_{E_{y}}(e_{x}(J/J^{2})e_{y},E_{y}).$
Proof.
See for example Proposition 2.4.3 in [3]. ∎
Now assume that the ext-quiver of $\operatorname{mod}A$ is acyclic. Then $\operatorname{End}_{A}(e_{x}A)\cong E_{x}$ for each $x$ by the above lemma. For any $A$-module $M$ of finite length, an induction on its length shows that the multiplicity of $S_{x}$ in a composition series is given by
$M:S_{x}]=\operatorname{dim}_{E_{x}}\operatorname{Hom}_{A}(e_{x}A,M).$
Lemma 7.
Let $A$ be a right artinian ring such that the ext-quiver of $\operatorname{mod}A$ is connected and acyclic. Then the following are equivalent.
  1. (1)
    For each simple $A$-module all its endomorphisms are central.
  2. (2)
    The centre $Z(A)$ of $A$ is a field and isomorphic to the endomorphism ring of each simple $A$-module.
Proof.
One direction is clear. Thus assume that (1) holds. We may identify $Z(A)=Z(\operatorname{mod}A)$. For any index $x$ let $I_{x}$ denote the kernel of the canonical homomorphism $Z(A)\to E_{x}$. Then $Z(A)/I\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}E_{x}$ since we assume that all elements of $E_{x}$ are central. For each pair $x,y$ the action of $Z(A)$ on
$e_{y}Ae_{x}\cong\operatorname{Hom}_{A}(e_{x}A,e_{y}A)$
factors through $E_{x}$ and through $E_{y}$, since the ext-quiver is acyclic. Thus $I_{x}=I_{y}$ when $e_{y}Ae_{x}\neq 0$. Because $A$ is connected it follows that $I_{x}=I_{y}$ for all $x,y$. Finally, $I_{x}=0$ for all $x$ since $I_{x}$ annihilates all elements of $A$. ∎
From now on assume that for each simple $A$-module all its endomorphisms are central and set $\mathbb{K}:=Z(A)$.
Lemma 8.
Let $M$ be an $A$-module such that $M:S_{x}]=1$ and $e_{x}A$ is a subquotient of $M^{r}$ for some $r\geq 1$. Then every non-zero morphism $e_{x}A\to M$ is a monomorphism.
Proof.
The assumption yields a monomorphism $\phi\colon e_{x}A\to M^{r}$ for some $r\geq 1$. Choose a non-zero morphism $\psi\colon e_{x}A\to M$. Then there are $\lambda_{i}\in E_{x}$ such that $\phi_{i}=\psi\lambda_{i}$ for $1\leq i\leq r$. We have
$\operatorname{Ker}\psi=\bigcap_{i}\operatorname{Ker}\psi\lambda_{i}=\operatorname{Ker}\phi=0$
where the first equality uses that the $\lambda_{i}$ are central. ∎
Now fix an $A$-module $M$ such that $M:S_{x}]=1$ for all $x$, and we assume that every non-zero morphism $e_{x}A\to M$ is a monomorphism. Then $e_{x}A$ is multiplicity free for all $x$ and there is a partial order on the set $P$ given by
$x\leq y\quad:\iff\quad e_{y}Ae_{x}\neq 0.$
Choose a $\mathbb{K}$-basis of $M$ that consists of elements $m_{x}\in Me_{x}$. This choice determines a $\mathbb{K}$-basis of $A$ consisting of elements $e_{xy}\in e_{y}Ae_{x}$ such that for all $x\leq y$
$m_{y}e_{xy}=m_{x}.$
For $x\leq y\leq z$ the equality
$m_{z}(e_{yz}e_{xy})=(m_{z}e_{yz})e_{xy}$
implies
(2) $e_{yz}e_{xy}=e_{xz}.$
We are ready to compare $A$ with the incidence algebra $\mathbb{K}P$ that is given by the poset coming from the ext-quiver of $\operatorname{mod}A$. I am grateful to Dieter Vossieck for suggesting a simplification of this argument.
Lemma 9.
The $\mathbb{K}$-linear map $A\to\mathbb{K}P$ given by the assignment $e_{xy}\mapsto a_{xy}$ for any $x\leq y$ is an isomorphism of $\mathbb{K}$-algebras.
Proof.
The map is an isomorphism of vector spaces; it is multiplicative because of the identities (1) and (2). ∎
Proof of the main theorem
Recall that the Hasse diagram of a poset $P$ is the quiver with vertex set $P$ and an arrow $x\to y$ if $x<y$ and there is no $z$ with $x<z<y$.
Proof of Theorem 1.
Let $\mathcal{A}$ be a length category. Suppose first that $\mathcal{A}$ is equivalent to $\operatorname{rep}(P,\mathbb{K})$ for some finite poset $P$ and a field $\mathbb{K}$. The simple representations identify with the elements of $P$. Specifically, for $x\in P$ the simple representation $S_{x}$ is given by $S_{x}(y)=\delta_{xy}\mathbb{K}$ for $y\in P$. In particular, it satisfies $\operatorname{End}(S_{x})\cong\mathbb{K}$. The ext-quiver identifies with the Hasse diagram of $P$. Thus there are no oriented cycles. Now consider the representation $M$ given by $M(x)=\mathbb{K}$ and $M(y)\to M(x)$ the identity map for all $x\leq y$. Then $M:S_{x}]=1$ for all $x$, and therefore the lattice of subobjects is distributive. The subrepresentation $M_{x}\subseteq M$ satisfying $M_{x}(y)\neq 0$ if and only if $y\leq x$ yields a projective cover of $S_{x}$. Thus every indecomposable projective representation embeds into $M$, and each object is a subquotient of $M^{r}$ for some $r\geq 1$.
Now suppose that conditions (1)–(3) hold for $\mathcal{A}$. It follows from Proposition 3 that $\mathcal{A}$ is equivalent to $\operatorname{mod}A$ for some Artin algebra $A$. More precisely, condition (2) implies that $\mathcal{A}$ is hom-finite; see Example 4. Then condition (3) implies that $\mathcal{A}$ admits a projective generator $Q$ such that $\operatorname{Hom}(Q,-)$ induces an equivalence $\mathcal{A}\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}A$ for $A:=\operatorname{End}(Q)$. Next observe that $\mathbb{K}:=Z(\mathcal{A})$ is a field by Lemma 7. Here we need to assume that $\mathcal{A}$ is connected. Using Lemmas 2 and 8, the existence of a distributive object implies that Lemma 9 can be applied. Thus $A\cong\mathbb{K}P$ for the poset $P$ given by the isomorphism classes of simple $A$-modules, and the equivalence with $\operatorname{rep}(P,\mathbb{K})$ follows from Lemma 5. ∎
3. Remarks
We conclude with several comments on the main theorem.
Remark 10.
Consider a length category $\mathcal{A}=\operatorname{rep}(P,\mathbb{K})$ as in Theorem 1. The distributive object $M$ arising in condition (3) is not necessarily unique. In fact, such objects are parameterised by 22-cocycles $\mu=(\mu_{xy})$ in $H^{*}(\Sigma(P),\mathbb{K}^{\times})$, where $\Sigma(P)$ denotes the simplicial complex associated with $P$ and for $x\leq y$ the map $M(y)\to M(x)$ sends the basis element $m_{y}$ to $m_{x}\mu_{xy}$; see [9, 10].
Remark 11.
In Theorem 1 condition (1) is needed. Let $\mathcal{A}=\operatorname{mod}\mathbb{K}[\varepsilon]$ where $\mathbb{K}[\varepsilon]$ denotes the algebra of duals numbers over a field. This is not a category of poset representations, though conditions (2) and (3) hold. Take for $M$ the indecomposable projective module.
The following example is interesting for several reasons, and I am grateful to Claus Michael Ringel for suggesting it. The example shows that in Theorem 1 condition (2) is needed. Also, it yields infinitely many objects $M$ of length two such that all objects are a subquotient of a finite direct sum of copies of $M$.
Example 12.
Consider the matrix algebra
$A=\begin{bmatrix}\mathbb{C}&\mathbb{C}\oplus\mathbb{C}_{\sigma}\\ 0&\mathbb{C}\end{bmatrix}$
where the action of $\mathbb{C}$ on $\mathbb{C}_{\sigma}$ from the right is twisted by conjugation; it is a finite dimensional $\mathbb{R}$-algebra. There are infinitely many isomorphism classes of length two $A$-modules, and all but two are faithful; see the Addendum of [5].
Remark 13.
In Theorem 1 the assumption on the length category $\mathcal{A}$ to be connected can be removed. Then one obtains a finite product $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ of connected length categories and possibly different fields $Z(\mathcal{A}_{1}),\ldots,Z(\mathcal{A}_{n})$.
Remark 14.
Any partially ordered set $P$ carries the structure of a topological space, by taking as open subsets the downward closed subsets of $P$. This assignment identifies finite posets with finite $T_{0}$-spaces [1, §1]. Moreover, for any abelian category $\mathcal{C}$ the category of sheaves on $P$ with values in $\mathcal{C}$ identifies with the category of representations $P^{\mathrm{op}}\to\mathcal{C}$, by taking a sheaf $\mathscr{F}$ on $P$ to the representation
$x\longmapsto\mathscr{F}_{x}=\mathscr{F}(\downarrow\!x)\qquad(x\in P)$
where
$\downarrow\!x:=\{y\in P\mid y\leq x\}.$
Thus Theorem 1 characterises the categories of sheaves on finite $T_{0}$-spaces with values in the category of finite dimensional vector spaces over a field. One may compare this with Gabriel’s reconstruction of a noetherian scheme from its category of quasi-coherent sheaves [7, VI].
Remark 15.
It would be interesting to have a version of Theorem 1 characterising categories of the form $\operatorname{rep}(P,\mathbb{K})$ when $\mathbb{K}$ is a skew field.
Acknowledgements
The author is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support. This work was supported by the Deutsche Forschungsgemeinschaft (SFB-TRR 358/1 2023 - 491392403).